OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (4, –6) and is parallel to y = –1/2x + 5. y = –2x + 2 y = 2x – 14 y = x – 8 y = –x – 4

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

it's simple .. where are you stuck ?

OpenStudy (anonymous):

on the question lol

OpenStudy (anonymous):

do you understand the question ?

OpenStudy (anonymous):

not at all

OpenStudy (anonymous):

i don't understand the equation and the parallel part

OpenStudy (anonymous):

\[y - y_{1} = m (x - x_{1})\] y - (-6) = (-1/2)(x - 4) y + 6 = -1/2x + 2 y = -1/2x - 4

OpenStudy (anonymous):

Do you understand what a straight line is ? and how it is represented ?

OpenStudy (anonymous):

The parallel line is a hint as to how to find the slope parallel lines have EQUAL slopes

OpenStudy (anonymous):

Any straight line $$ax+by=c$$ is parallel to $$ax+by=k$$

OpenStudy (anonymous):

wow i couls never get this graph stuff no matter how hard i try